How many milliliters of 0.100 M NaOH are required to neutralize 65.0 mL of 0.250 M H2SO4 ? The balanced neutralization reaction is:
H2SO4(aq)+2NaOH(aq)→Na2SO4(aq)+2H2O(l).
Answer #1

mmol of acid = MV = 65*0.25 = 16.25 mmol of H2SO4 = 16.25*2 H+ = 32.5 mmol of H+

then

mmol of OH- = 32.5 mmol

V = mmol/M = 32.5/(0.1) = 325 mL of base needed